OpenStudy (anonymous):

soybean meal is 18% protein cornmeal is 9% protein how many pounds of each should be mixed together in order to get 360 lbs mixture that is 11% protien.... Help please

OpenStudy (anonymous):

Call each kind by a variable: Let x be the amount of Soybean y ---------------- cornmeal The total amount of both kinds: x + y = 360 The mixture: 0.18x + 0.09y = 0.11(360) = 39.6 So you solve the system of two linear equations: x + y = 360 0.18x +0.09y = 39.6 Hint: 80 lbs of SOYBEAN and 280 lbs of CORNMEAL

OpenStudy (anonymous):

can you solve?

OpenStudy (anonymous):

x + y = 360 y= 360-x subtitution : y= 360-x 0.18x +0.09*(360-x) = 39.6 0.18x-0.09x = 39.6 - 32.4 0.09x = 7.2 x = 7.2 : 0.09 x = 80 Subtitution x = 80 , find y x + y = 360 80+y= 360 y = 360 -80 y = 280 80 lbs of SOYBEAN and 280 lbs of CORNMEAL